Compute the answers to the questions below with the given interest rate.

Interest rate (annual) 6%
Year Cash Flow
1 $            4,257
2 $            4,257
3 $            4,257
4 $            4,257
5 $            4,257
6 $            4,257
Annuity Due Present Value
Annuity Due Future Value
Ordinary Annuity Present Value
Ordinary Annuity Future Value
Answer #1

Solution:

Annuity due present value = $4,257 * Cumulative PV factor at 6% for 6 periods of annuity due

= $4,257 * 5.21236 = $22,189

Annuity due future value = $4,257 * Cumulative FV factor at 6% for 6 periods of annuity due

= $4,257 * 7.39384 = $31,476

Ordinary annuity present value = $4,257 * Cumulative PV factor at 6% for 6 periods

= $4,257 * 4.91732 = $20,933

Ordinary annuity future value = $4,257 * cumulative FV factor at 6% for 6 periods

= $4,257 * 6.97532 = $29,694